OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the point (5, −3) and has a slope of −1.25. y=___x+___

pooja195 (pooja195):

ok well you know the point slope formula right?

pooja195 (pooja195):

y-y1=m(x-x1) right?

pooja195 (pooja195):

using this formula plug in the numbers

pooja195 (pooja195):

y+3=-1.25(x-5) solve it

pooja195 (pooja195):

@redflame12 can you atleast try?

pooja195 (pooja195):

ok first u always distribute so 1.25xX=? then 1.25x-5=?

pooja195 (pooja195):

ok u get y+3=1.25x-6.25

pooja195 (pooja195):

then u add 3 to both sides and your answer would be y=1.25x-3.25